OpenStudy (anonymous):

Solve (no decimals): -3x( x -7)(-4x +11)= 0

OpenStudy (anonymous):

is this a linear equation?

hero (hero):

Hint: -3x = 0 x - 7 = 0 -4x + 11 = 0

OpenStudy (anonymous):

so...the whole thing =0?

hero (hero):

You have to solve each equation for x.

hero (hero):

@ravenbypoe

OpenStudy (anonymous):

mmm .. as you can see, I am just not good in math. do I subtract each number on each side to get the variable alone?

hero (hero):

The goal is to isolate x in each equation. In order to do that you must perform inverse operations. For example, if a number is multiplied by x, you divide by x on both sides of the equal sign to isolate x. If a number is added to x, then you subtract if from x on both sides.

hero (hero):

For example 3x + 6 = 0 Subtract 6 from both sides to get 3x = -6 Then divide both sides by 3 to get x = -6/3 Simpify that to get x = -2

hero (hero):

@ravenbypoe

OpenStudy (anonymous):

thank you

hero (hero):

Have you solved the equations yet? That what I demonstrated above is only an example.

OpenStudy (anonymous):

got the example, makes sense, but in this the format is not 3x+6...is -3x=0 so im not sure...i tried to do the multiplication of -3x by the 2 parenthesis but then i dont know what else. thank you for your help ;) so i got -3x^2+21-4x^2-11x+28x-77... but this cant be right....

hero (hero):

Yes, you need plenty help with this. You were asked to solve for x, not re-multiply the whole thing.

OpenStudy (anonymous):

yep-i know...:(